Let

f

be a differentiable function on

[Alpha, beta ]

and

x belongs to [alpha, beta ].

Show that, if

f '(x) = 0

and

f ''(x) =0,

then

f

must have a local maximum at

x.


1
Expert's answer
2021-05-07T09:30:49-0400

As "x\in [\alpha,\beta]"


also f is a differentiable function


According to mean value theorem-


"f'(x)=\dfrac{f(\beta)-f(\alpha)}{\beta-\alpha}"


"f(\beta)-f(\alpha)=0\Rightarrow f(\beta)=f(\alpha)"


As f'(x) is 0 then, ,

There exist a critical point between "[\alpha,\beta]"


The function does not calls us nothing about the maximum or minimum.


As f''(x) =0, The given test does not tells nothings.


f(X) has local maximum at x.